# Exploring filament galaxies using AstroSat/ UVIT

### v1

@@ Abstract
We present results from our deep Far-ultraviolet (FUV) survey using AstroSat/UVIT of a filamentary structure at . A total of four filaments comprising 58 galaxies were probed in our study. We detect 18 filament galaxies in our FUV observation. All filament galaxies are further classified based on their photometric color, nuclear activity, and morphology. The filaments contain galaxies with mixed stellar population types and structures. We do not detect galaxies in our UVIT survey up to a distance of 0.4 Mpc from the filament axis, implying a dearth of recent star formation in the inner region of filaments. We witness an increase in FUV specific star formation rate (sSFR) of filament galaxies with increasing distance from the filament spine (D). On the contrary, no such gradient in FUV sSFR with cluster-centric distance is observed in the case of cluster galaxies. The high stellar mass filament galaxies (M 10 M) were more star-forming than cluster galaxies in a fixed mass range. The FUV morphology of some filament galaxies detected in the filament outskirts (D 0.9 Mpc) is comparable to or slightly extended than their optical counterpart. The mass assembly of galaxies probed by estimating color gradients show that more centrally star-forming galaxies reside closer to the filament axis regardless of stellar mass. Our results prove that the likelihood of merger interaction and gas starvation increases on approaching the filament spine. We report a definitive and in-homogeneous impact of filaments on the galaxies residing inside them.
@@ 1 Introduction
Filaments are the most visually dominant component of the cosmic web. Around 50% of galaxies or baryonic mass in the Universe resides within filaments while the structure occupies only 5% of the cosmic volume (Cautun et al., 2014). Filaments inhabit the spaces between massive clusters and inside voids (Tempel et al., 2014). Filaments are not just an assembly of galaxies; they provide a dynamical pathway in channeling the galaxies to massive clusters. Apart from the cool baryonic matter in galaxies, filaments are shown to be the carrier of warm-hot intergalactic matter (WHIM; Galárraga-Espinosa et al., 2021). The filamentary pattern in the cosmic web is well-illustrated both in simulations (Colberg et al., 2005; Cautun et al., 2014) as well as observations (e.g., Durret et al., 2003; Porter & Raychaudhury, 2005). The detection of filaments within the cosmic web is an intricate task due to their low environmental density. There are various algorithms available for extracting filaments in the cosmic web, such as Bisous (Tempel et al., 2014), DisPerSE (Sousbie et al., 2011), NEXUS+ (Cautun et al., 2013, 2014), etc.
@@
Several large-scale phenomena, such as cosmic web enhancement and detachment, are reported to occur inside filaments. A few filament galaxies are shown to host ionized gas clouds in their circumgalactic region due to cosmic web enhancement (Vulcani et al., 2019). On the other hand, it is theoretically indicated that galaxies experience gas starvation caused by cosmic web detachment inside filaments due to the non-linear interaction of the cosmic web (Aragon Calvo et al., 2019). These physical mechanisms affect the ongoing star formation activity of the constituent galaxies. Filaments also affect their resident galaxies dynamically due to non-uniform large-scale matter distribution (Lee & Pen, 2000; Tempel & Libeskind, 2013). Therefore, it is evident that the filaments significantly impact the evolution of the galaxies inside them.
@@
There are a number of observation- (e.g., Kuutma et al., 2017; Mahajan et al., 2018; Lee et al., 2021) and simulation-based (e.g., Singh et al., 2020) studies performed to probe the structural and physical properties of filament galaxies. Most of these analyses are based on optical imaging and spectroscopic datasets. The local environment density around filament galaxies and the distance of galaxies from the filament spine are found to impact their growth (Bonjean et al., 2020). Studies focusing on undermining the properties of galaxies as a function of their distances from the spine of filaments found that the passive galaxy fraction increases closer to the filament axis (Kuutma et al., 2017; Laigle et al., 2018; Singh et al., 2020; Bonjean et al., 2020). The authors report a morphological transformation of filament galaxies from late- to early-type, followed by a suppression in ongoing star formation when approaching the filament spine from the void environment. An increased likelihood of finding bright and red galaxies at a similar large-scale luminosity density is observed in filaments than outside them due to a higher number of elliptical galaxy populations (Poudel et al., 2017). Although filaments are found to host red and passively evolving galaxies, they are comparatively less massive and bluer than their counterpart in the clusters (Mahajan et al., 2018). Meanwhile, filaments are reported to act as a stimulator for star formation activities in galaxies, e.g., Darvish et al. (2014) uncovers the possibility of enhanced gas accretion in filament galaxies through inter-filamentary channels. Filaments provide an active environment to trap cold gas, assist its cooling, and enhance star formation activity (Liao & Gao, 2019). The discussion above reveals an inhomogeneity in the star formation properties among filament galaxies.
@@
The ultraviolet (UV) emission broadly divided into two wavelength ranges, far-UV (FUV) and near-UV (NUV), from galaxies is an ideal tracer of recent star formation in galaxies (Gil de Paz et al., 2007). A few investigations in the past have probed the properties of filament galaxies with the help of UV and UV-optical colors (e.g., Mahajan et al., 2018; Lee et al., 2021) using observations from Galaxy Evolution Explorer (GALEX, Morrissey et al., 2007). The authors examine the gradient in UV colors of filament galaxies when approaching the filament spine. Another study inspected the changes in GALEX-based NUV star formation rates (SFRs) and stellar mass of spiral galaxies along filaments (Alpaslan et al., 2016). These analyses hinted towards a definitive transformation in UV properties along the filament. However, none of the studies have focused on deep FUV properties of filament galaxies as GALEX All-sky Imaging (AIS) survey, which covers a maximum portion of the sky and has a typical exposure time of 100 s with a resolution of 5 (Morrissey et al., 2007). The FUV emission provides a clear separation between recently star-forming and passive galaxies. The average timescale of FUV emission is roughly 0.1 Gyr, even shorter than the NUV emission timescale of about 1 Gyr (Gil de Paz et al., 2007). Therefore, deep and spatially-resolved FUV observation of filaments could help us trace the ongoing recent star formation along the filaments.
@@
This work is based on a proposed deep FUV survey of filamentary structure conducted by AstroSat/ Ultraviolet Imaging Telescope (UVIT, Tandon et al., 2017). The rectangular box shown in Figure 1 encloses the targeted filamentary structure. An enhanced spatial resolution ( three times better than GALEX) and greater sensitivity of UVIT would help us identify recently star-forming galaxies in the filaments, which may be undetected in previously done shallow UV imaging surveys. We have conducted a similar deep UVIT survey of the Bootes void (red circle in Figure 1; Kirshner et al., 1981). The observations led to the detection of a few void galaxy candidates, which were undetected in GALEX observation. We reported a distinct blueward shift in color of void galaxies compared to a sample of nearby galaxy (Pandey et al., 2021). Voids and clusters are the rarest and densest environments, respectively. At the same time, filaments are thought to represent the intermediate environment density. A detailed understanding of star formation activity across environments will be possible when enough studies explore these environments in FUV. Our FUV observation of the filaments would be utilized to investigate the gradient in UV emission across its width. We plan to compare the results with galaxies in different environments.
Physical phenomena such as galaxy mergers, gas starvation and accretion, etc., occur inside filaments. These mechanisms, along with the star formation history of a galaxy, can lead to different stellar mass assembly mechanisms (Lin et al., 2019). We would compare the FUV morphology of filament galaxies with their optical to identify mergers or gas accretion signatures. Moreover, we plan to utilize multi-wavelength observations to probe the mass assembly of filament galaxies to highlight the dominant physical mechanisms responsible for galaxy evolution in filaments. This study would provide insights into the role of cosmic filaments in shaping the star formation activity of galaxy populations.
@@
This paper is organized as follows: Section 2 describes the observational data used in the work and UVIT data reduction techniques followed, and Section 3 explains our methodology applied in extracting filament galaxies from UVIT observations. We derive the photometric and chemical properties of filament galaxies in Section 4. Various aspects of FUV emission of the sample, such as morphology and FUV SFRs are discussed in Section 5. In Section 6, we examine the mass assembly of filament galaxies. Our findings are discussed in Section 7. A standard CDM cosmology with = 0.3, = 0.7, and H = 70 km s Mpc is assumed in this paper.
@@ 2.1 Observations
We primarily use proposed AstroSat/UVIT data observed in BaF2 filter centered at Å (Observation ID: A09_127T01_9000003816, A09_127T03_9000003820; PI: Divya Pandey). The total field-of-view (FOV) of the UVIT channel is 28 in diameter. The pixel scale of the reduced image 0.417, whereas the zero-point of the filter is 17.78 mag (Tandon et al., 2020). Under the proposal, we observe a filamentary structure at z 0.073. The observed structure is located close to one of the largest voids in the northern hemisphere, the Bootes void (See Figure 1). We proposed to observe the structure in two FOVs with an exposure time of 7 kilo seconds each. The central coordinates of the two pointing are: = 217.15, = 46.12 (FOV2) and = 215.85, = 46.1 (FOV1). Figure 1 shows the direction of UVIT pointings on the cosmic web.
@@ 2.2 UVIT data reduction and analysis
We obtain the raw Level 1 version of the observations from AstroSat data archive111https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp. The level 1 data is fed into CCDLAB (Postma & Leahy, 2021). CCDLAB is an automated pipeline used for reducing the L1 version of UVIT data into science-ready images. The data reduction includes drift correction, flat-field correction and is further corrected for additional distortion and fixed noise pattern. The corrected images from different orbits are stacked together through the centroid alignment process to obtain a deep, science-ready image. CCDLAB removes frames containing large numbers of photon counts due to cosmic ray shower, leading to a reduction in the total exposure time of the image. The effective exposure time (t) of the final images was 5.7 ks and 6.3 ks. The world coordinate system (WCS) was assigned to the final images using Astroquery (Ginsburg et al., 2019) appended with CCDLAB. Astroquery uses the GAIA (Gaia Collaboration et al., 2016) catalog to assign coordinates to each pixel in the final image. The mean error in the astrometry was 0.28. Along with the science-ready images, we get corresponding exposure maps as an output from the CCDLAB.
We run source extractor (Bertin & Arnouts, 1996) for detecting sources observed in the two sets of FUV images. The background is detected in automated mode. The detection and analysis threshold is fixed at 1.5 for source extraction and photometry. We consider the minimum number of pixels required to classify an intensity distribution as a source (DETECT_MINAREA) to be 9. The 3 limiting magnitudes for our observation in FOV 1 and FOV 2 are 24.79 mag and 25.07 mag, respectively. We detected a total of 3574 sources combining both the FOVs. We perform Petrosian photometry to measure the integrated magnitude of the detected sources. We expect to detect irregular and disky galaxies in our UVIT FOVs wherein the Petrosian apertures may recover nearly all flux from an exponential light distribution (Blanton et al., 2001).
@@ 3 Identifying filament galaxies
Only a fraction of the detected sources in our the FUV observation belong to the filaments. Hence, the identification of filament galaxies is imperative part of our study. Multiple catalogs employ various algorithms to find filaments and associated galaxies in large-scale structures (Tempel et al., 2014; Carrón Duque et al., 2022; Alpaslan et al., 2014). We use a catalog given by Tempel et al. (2014, hereafter Tempel catalog) in our study. The authors use the Bious model to detect filaments in an extensive database of galaxies observed with SDSS. The catalog provides a detailed list of filaments and constituent filament galaxies. The information regarding the distance from the nearest filament spine (D), comoving distance, integrated magnitudes, and optical luminosities for each filament galaxy are included in the catalog.
We matched our UVIT source catalog with the Tempel catalog within a radius of 2.5 to identify filament galaxies in our FUV observations. The detected galaxies belong to four different filaments (Filament ID: 1029, 4188, 9987, and 9374 in the Tempel catalog) of different lengths. The four filaments together contain 58 galaxies, of which 18 were detected in the UVIT sample. We restrict our filament sample up to D = 1 Mpc, as previous reports on filament galaxies point out that the filament considerably impacts the properties of galaxies situated within 1 Mpc distance from their axis (Tempel & Tamm, 2015; Tempel et al., 2015). A sky map depicting the distribution of galaxies in each filament is shown in Figure 2. The inset figure highlights the UVIT FOVs, UVIT targeted and detected galaxies. The red-colored markers within FOVs represent filament galaxies undetected in our survey, whereas the galaxies depicted by blue markers were detected in the FUV survey. Prior to UVIT, a shallow FUV survey (t 100 s) of the observed FOVs was conducted by GALEX. We find that six of the 18 filament galaxies detected by UVIT remained missing in GALEX AIS catalog. In other words, these filament galaxies are newly detected in our UVIT survey. Table 1 consists of information on each filament, their constituent galaxies, and a comparison between filament galaxies detected by UVIT and GALEX AIS observation.
| Fil. ID | Length | Total Galaxies | FUV | FUV |

### v2

@@ Abstract
We present results from our deep Far-ultraviolet (FUV) survey using AstroSat/UVIT of a filamentary structure at . A total of four filaments comprising 58 galaxies were probed in our study. We detect 18 filament galaxies in our FUV observation. All filament galaxies are further classified based on their photometric color, nuclear activity, and morphology. The filaments contain galaxies with mixed stellar population types and structures. We do not detect galaxies in our UVIT survey up to a distance of 0.4 Mpc from the filament axis, implying a lack of recent star formation in the inner region of filaments. The FUV star formation rate (SFR) for star-forming galaxies agrees well with the SFR144MHz calculated using Low-Frequency Array (LOFAR) radio-continuum observations. We witness an increase in FUV specific-SFR (sSFR) of filament galaxies with increasing distance from the filament spine (Dfil). The intermediate-to-high stellar mass filament galaxies were more star-forming than cluster galaxies in a fixed stellar mass bin. The FUV morphology of some filament galaxies detected in the filament outskirts (Dfil 0.7 Mpc ) is comparable to or slightly extended than their optical counterpart. The mass assembly of galaxies examined by estimating color gradients show that more ‘red-cored’ galaxies reside in the outer region of the filaments. Our results prove that the likelihood of merger interaction and gas starvation increases when approaching the filament spine. We report a definitive and in-homogeneous impact of filaments on the galaxies residing inside them.
@@ 1 Introduction
Early redshift surveys (e.g. de Lapparent et al., 1986; Colless et al., 2001) and numerical simulations (e.g. Efstathiou et al., 1985) reveal the large-scale structure (LSS), comprising clusters, voids, walls and filaments (Bond et al., 1996). The filaments are the most dominant LSS as around 50% of the total mass fraction in the cosmic web resides within filaments, while the LSS occupies only 5% of the cosmic volume (Cautun et al., 2014). Filaments inhabit the spaces between massive clusters and inside voids (Tempel et al., 2014). The LSS is not just an assembly of galaxies; it provides a dynamic pathway in channeling the galaxies to massive clusters. Apart from the cool baryonic matter in galaxies, filaments are shown to be the carrier of warm-hot intergalactic matter (WHIM; Galárraga-Espinosa et al., 2021). The filamentary pattern in the cosmic web is well-illustrated both in simulations (Colberg et al., 2005; Cautun et al., 2014) as well as observations (e.g., Durret et al., 2003; Porter & Raychaudhury, 2005). The detection of filaments within the cosmic web is an intricate task due to their variable thickness depending on the surrounding environment density. There are various algorithms available for extracting filaments in the cosmic web, such as Bisous (Tempel et al., 2014), DisPerSE (Sousbie et al., 2011), NEXUS+ (Cautun et al., 2013, 2014).
@@
Several large-scale phenomena, such as cosmic web enhancement and detachment, are reported to occur inside filaments. A few filament galaxies are shown to host ionized gas clouds in their circumgalactic region due to cosmic web enhancement (Vulcani et al., 2019). On the other hand, it is theoretically indicated that galaxies experience gas starvation caused by cosmic web detachment inside filaments due to the non-linear interaction of the cosmic web (Aragon Calvo et al., 2019). The large-scale matter distribution along filament is found to be highly non-uniform (Lee & Pen, 2000; Tempel & Libeskind, 2013; Cautun et al., 2014), depending on where the filaments originated. Detailed study of the cosmic structure shows that the thickness of filaments, as well as the angle between galaxy halo spin and filament spine, profoundly impact the properties of resident galaxies (Ganeshaiah Veena et al., 2018, 2019, 2021).
@@
There are several observation- (e.g., Tempel & Libeskind, 2013; Kuutma et al., 2017; Mahajan et al., 2018; Lee et al., 2021) and simulation-based (e.g., Ganeshaiah Veena et al., 2018, 2019; Singh et al., 2020; Hasan et al., 2023) studies performed to probe the structural and physical properties of filament galaxies. Most of these analyses are based on optical imaging and spectroscopic datasets. A few factors, such as the local environment density, the distance of galaxies from the filament spine, and the angle between galaxy halo spin alignment and filament axis, impact the galaxy evolution inside filaments (Bonjean et al., 2020; Ganeshaiah Veena et al., 2018, 2021). Studies focusing on characterizing the properties of galaxies as a function of their distances from the spine of filaments found that the passive galaxy fraction increases closer to the filament axis (Kuutma et al., 2017; Laigle et al., 2018; Singh et al., 2020; Bonjean et al., 2020). The authors report a morphological transformation of filament galaxies from late- to early-type, followed by a suppression in ongoing star formation when approaching the filament spine from the void environment.
The spin alignment of galaxies in filaments is another determinant of the properties of galaxies. Tempel & Libeskind (2013) presents strong observational evidence suggesting that the spins of spiral galaxies align parallel to the filament axis, whereas the minor axis of elliptical galaxies typically orient themselves perpendicular to the filament axis. Ganeshaiah Veena et al. (2018, 2019) highlight mass dependence on the halo spin and filament alignment. The authors show that low-mass halos tend to reside parallel to the filament axis and high-mass halos align preferentially perpendicular to the filament. These observations are a manifestation of large-scale tidal fields explained by the tidal torque theory (TTT; Hoyle et al., 1949; Peebles, 1969; White, 1984). The tidal fields affect gas inflow history inside filaments, impacting the ongoing star formation.
@@
The ultraviolet (UV) emission broadly divided into two wavelength ranges, far-UV (FUV) and near-UV (NUV), from galaxies is an ideal tracer of recent star formation in galaxies (Gil de Paz et al., 2007). The average timescale of FUV emission is roughly 0.1 Gyr, even shorter than the NUV emission timescale of about one Gyr. A few investigations in the past have probed the properties of filament galaxies with the help of UV and UV-optical colors (e.g., Alpaslan et al., 2016; Mahajan et al., 2018; Lee et al., 2021). The authors examine the gradient in UV colors, NUV star formation rates (SFRs), and stellar mass of filament galaxies when approaching the filament spine. These analyses hinted towards a definitive transformation in UV properties along the filament. Most of these investigations are based on the observations from Galaxy Evolution Explorer ( GALEX, Morrissey et al., 2007) which has performed a shallow survey of the entire sky having a typical exposure time of 100 s with a resolution of 5′′ (Morrissey et al., 2007). It is worth noting that a deep and resolved FUV survey of the filament environment would be helpful to trace star formation activity in filaments even at the faint magnitude limit.
@@
This work is based on a proposed deep FUV survey of filamentary structure conducted by AstroSat/ Ultraviolet Imaging Telescope (UVIT, Tandon et al., 2017). The rectangular box in Figure 1 encloses the targeted filamentary structure ( 0.07). The UVIT provides high spatial resolution ( three times better than GALEX) and enhanced sensitivity that would be utilized to identify recently star-forming galaxies inside filaments, which would otherwise remain undetected in previous UV imaging surveys. We have conducted a similar deep UVIT survey of the Bootes void (red circle in Figure 1; Kirshner et al., 1981). The observations led to the detection of a few void galaxy candidates, which were not identified in GALEX survey. We reported a distinct blueward shift in color of void galaxies compared to a sample of local galaxies (Pandey et al., 2021). Unlike voids and galaxy clusters, representing the rarest and densest environment in the cosmic web, the filaments host intermediate environment density. A detailed understanding of star formation activity across different LSSs will be possible when enough studies explore these environments in FUV.
Substantial evidence has been reported in support of amplification of the star formation activity in galaxies residing in filaments, e.g., Darvish et al. (2014) uncovers the possibility of mild galaxy-galaxy interaction or efficient gas accretion through inter-filamentary channels in filament galaxies. Filaments provide an active environment to trap cold gas, assist its cooling, and enhance star formation activity (Liao & Gao, 2019). Utilizing the deep UVIT observations of filaments, we compute SFRs and compare the FUV and optical morphology of filament galaxies to uncover signs of interactions or gas accretion signatures. We also probe the stellar mass-assembly history of filament galaxies using FUVoptical colors in this work. The results are compared with galaxies residing in different LSSs.
@@
This paper is organized as follows: Section 2 describes the observational data used in the work and UVIT data reduction techniques followed, and Section 3 explains our methodology applied in extracting filament galaxies from UVIT observations. We derive the photometric and chemical properties of filament galaxies in Section 4. Various aspects of FUV emission of the sample, such as morphology and FUV SFRs are discussed in Section 5. In Section 6, we examine the mass assembly of filament galaxies. Our findings are discussed in Section 7. A standard CDM cosmology with = 0.3, = 0.7, and H0 = 70 km s-1 Mpc-1 is assumed in this paper.
@@ 2.1 Observations
We primarily use proposed AstroSat/UVIT data observed in BaF2 filter centered at Å (Observation ID: A09_127T01_9000003816, A09_127T03_9000003820; PI: Divya Pandey). The total field-of-view (FOV) of the UVIT channel is 28′ in diameter. The pixel scale of the reduced image is 0.417′′, whereas the zero-point of the filter is 17.78 mag (Tandon et al., 2020). Under the proposal, we observe a filamentary structure at 0.073. The observed structure is located close to one of the largest voids in the northern hemisphere, the Bootes void (See Figure 1). We proposed to observe the structure in two FOVs with an exposure time of 7 kilo seconds each. The central coordinates of the two pointing are: = 215.85∘, = 46.1∘ (FOV1) and = 217.15∘, = 46.12∘ (FOV2). Figure 1 shows the direction of UVIT pointings on the cosmic web.
@@ 2.2 UVIT data reduction and analysis
We obtain the raw Level 1 version of the observations from AstroSat data archive111https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp. The level 1 data is fed into CCDLAB (Postma & Leahy, 2021). CCDLAB is an automated pipeline used for reducing the L1 version of UVIT data into science-ready images. The data reduction includes drift correction, flat-field correction and is further corrected for additional distortion and fixed noise pattern. The corrected images from different orbits are stacked together through the centroid alignment process to obtain a deep, science-ready image. CCDLAB removes frames containing large numbers of photon counts due to cosmic ray shower, leading to a reduction in the total exposure time of the image. The effective exposure time (texp) of the final images was 5.7 kilo seconds and 6.3 kilo seconds. The world coordinate system (WCS) was assigned to the final images using Astroquery (Ginsburg et al., 2019) appended with CCDLAB. Astroquery uses the GAIA (Gaia Collaboration et al., 2016) catalog to assign coordinates to each pixel in the final image. The mean error in the astrometry was 0.′′28. Along with the science-ready images, we get corresponding exposure maps as an output from the CCDLAB.
We run source extractor (Bertin & Arnouts, 1996) for detecting sources observed in the two sets of FUV images. The background is detected in automated mode. The detection and analysis threshold is fixed at 1.5 for source extraction and photometry. We consider the minimum number of pixels required to classify an intensity distribution as a source (DETECT_MINAREA) to be 9. The 3 limiting magnitudes for our observation in FOV 1 and FOV 2 are 24.79 mag and 25.07 mag, respectively. We detected a total of 3574 sources combining both the FOVs. We perform Petrosian photometry to measure the integrated magnitude of the detected sources. We expect to detect irregular and disky galaxies in our UVIT FOVs where the Petrosian apertures may recover nearly all flux from an exponential light distribution (Blanton et al., 2001).
@@ 3 Identifying filament galaxies
Only a fraction of the detected sources in our FUV observation belong to the filaments. Hence, identifying filament galaxies is an imperative part of our study. Multiple catalogs employ various algorithms to find filaments and associated galaxies in the LSS (Tempel et al., 2014; Carrón Duque et al., 2022; Alpaslan et al., 2014). We use a catalog given by Tempel et al. (2014, hereafter Tempel catalog) in our study. Tempel et al. (2014) employs the Bisous model to detect filaments in an extensive database of galaxies observed with SDSS.
Bisous algorithm randomly distributes fixed-sized cylinders on a large-scale galaxy distribution. The probability of a cylinder corresponding to a filamentary structure is based on the galaxy density contrast inside and outside the cylinder. Multiple Markov Chain Monte Carlo (MCMC) simulations are run to establish a network of well-aligned cylinders of fixed radius, varying length, and orientation. Using MCMC simulation, a visit map is created, which gives the probability of a region belonging to the filamentary structure. The ridges of this visit map where the galaxy density is highest are considered as the filament spine. Tempel Catalog provides a detailed list of filaments and constituent filament galaxies. The information regarding the distance from the nearest filament spine (Dfil), comoving distance, integrated magnitudes, and optical luminosities for each filament galaxy are included in the catalog. Tempel et al. (2014) apply density-dependent peculiar velocity sampling scheme (Kitaura & Enßlin, 2008) to mitigate finger-of-god effect (Tully & Fisher, 1978) in the galaxy distances used for filament extraction. The projection effect may also impact the distances measured in this study. As a result, the Dfil parameter used in this study may represent the lower limit of the actual separation between galaxies and filament spine.
We matched our UVIT source catalog with the Tempel catalog within a radius of 2.′′5 to identify filament galaxies in our FUV observations. The detected galaxies belong to four different filaments (Filament ID: 1029, 4188, 9987, and 9374 in the Tempel catalog) of different lengths. The four filaments together contain 58 galaxies, of which 18 were detected in the UVIT sample. We restrict our filament sample up to a distance of 1 Mpc from the filament spine (Dfil), as previous report points that filaments considerably impact the properties of galaxies situated within 1 Mpc from their axis (Kuutma et al., 2017). A sky map depicting the distribution of galaxies in each filament and corresponding filament spine is shown in Figure 2. The inset panels mark the position of UVIT pointings. Filament galaxies detected above the defined threshold are outlined with blue crosses, whereas those undetected are highlighted with red crosses. Prior to UVIT, a shallow FUV survey (texp 100 s) of the observed FOVs was conducted by GALEX. We find that six of the 18 filament galaxies detected by UVIT remained missing in the GALEX AIS catalog. In other words, these filament galaxies are newly detected in our UVIT survey. Table 1 consists of information on each filament, their constituent galaxies, and a comparison between filament galaxies detected by UVIT and GALEX AIS observation.
| Fil. ID | Length | Total Galaxies | FUVUVIT | FUVGALEX |